OpenStudy (mendicant_bias):

(Fourier Transform) I'm trying to solve an integral equation using Fourier Transforms, but the thing that I'm not used to/am not sure how to grapple with is that the integral equation itself is piecewise-defined; should I put it in the form of a linear combination of Heaviside Functions or something, or how do I deal with this? Original prompt below shortly.

OpenStudy (mendicant_bias):

\[\text{Solve the \integral equation}\] \[ \int\limits_{0}^{\infty}f(t)\sin(\alpha t)dt = \left\{ \begin{array}{lr} -\alpha, & 0<\alpha<1\\ 0, & 1<\alpha< \infty \end{array} \right. \]

OpenStudy (mendicant_bias):

@dan815

OpenStudy (amistre64):

how would you work this if it wasnt pieced?

OpenStudy (mendicant_bias):

I'd take the appropriate fourier transforms, or maybe inverse fourier transforms, I guess. I just realized that the LHS is identical to a Fourier Sine transform, so I guess I could take the inverse of both sides. That would give me the LHS just being the original function, so that's good, but I have no idea what the RHS would be.

OpenStudy (amistre64):

are you comfortable with switches?

OpenStudy (amistre64):

\[\int f=\mu_{(0,1)}(-\alpha)\] now when we get to 1 and onwards, switch 2 turns on so that we add to 0 \[\int f=\mu_{(0,1)}(-\alpha)+\mu_{(1,\infty)}(\alpha)\]

OpenStudy (mendicant_bias):

e.g. Heaviside functions, right? I've never seen that notation nor heard them called switches, but that's what it sounds like, which is what I mentioned in my original post.

OpenStudy (dan815):